How many ounces of a 16% alcohol solution and a 42% alcohol solution must be combined to obtain 78% ounces of a 24% solution.
Answer by Boreal(15235)

the ounces of the 16% are x
the ounces of the 42% are everything else, which is 78-x
.16x+.42(78-x)=78*.24, the amount of "pure" alcohol
.16x+32.76-.42x=18.72
-.26x=-14.04
x=54 ounces of 16%
78-x=24 ounces of 42%